# ManiFeel: Benchmarking and Understanding Visuotactile Manipulation Policy Learning

### v1

@@ F.3 Runtime Analysis of Tactile Representations
In this section, we evaluate the training time required for different tactile encoders and representations, including ResNet18, UniT, T3, and AnyTouch. Table 8 summarizes the training time per epoch and total training time (over 20 epochs) for visuotactile policies in the real-world sorting task under normal lighting. All experiments are conducted using an NVIDIA A30 GPU with a batch size of 8. AnyTouch incurs substantially higher computational cost, with each training epoch taking over 67 minutes. In contrast, UniT and T3 are more efficient, with training times of approximately 26 and 18 minutes per epoch, respectively, making them better suited for faster policy iteration and development.
@@
Table 8: Average training time per epoch and total training time (over 20 epochs) for visuotactile policies with different tactile representations in the real-world sorting task under normal lighting conditions. The dataset includes 103 human demonstrations. Time per epoch is reported in minutes and total time in hours.
@@
F.4 Experiment on Different Camera Viewpoints
@@
(b) Training Curve
FrontWrist005050505010010010010042.242.242.242.284.884.884.884.855.455.455.455.492.792.792.792.7Success rate (%)Vision-onlyVisuotactile
(c) Success Rate
Figure 7: Comparison of front and wrist camera views for vision-only and visuotactile policies in the tactile exploration task. (a) Sample observations showing tactile inputs and visual images from occluded front views and dim wrist views. (b) Training curves comparing policy performance across different viewpoints. (c) Final success rates (%) averaged over the last ten epochs and three seeds.
Figure 8: Rollout of the tactile sorting task in the simulation environment under dim lighting conditions. The policy takes front-view and tactile images as multimodal inputs. Note that tactile images are resized to a square shape of $256{\times}256$, from their original resolution of $240{\times}320$.
Figure 9: Rollout of the blind insertion task in the simulation environment. The policy takes front-view and tactile images as multimodal inputs. The agent view is shown for visualization purposes only and is not part of the policy’s observation inputs. Note that tactile images are resized to a square shape of $256{\times}256$, from their original resolution of $240{\times}320$.
Fig. 7 illustrates the performance of policies in the tactile exploration task when using visual observations from different camera viewpoints, specifically wrist-mounted and front-facing perspectives. Fig. 7(a) illustrate tactile inputs and visual images from occluded front views and dim wrist views. The results (Figs. 7(b)7(c)) show that both vision-only and visuotactile policies achieve higher success rates when trained with wrist-mounted inputs. Moreover, the visuotactile policy consistently outperforms its vision-only counterpart under both viewpoints, demonstrating that incorporating tactile sensing enhances performance under visually limited conditions.
(a) Normal lighting conditions.
(b) Dim lighting conditions.
Figure 10: Rollouts of the visuotactile policy in the real-world tactile sorting task under (a) normal and (b) dim lighting conditions. The policy receives observations from a front-facing camera and a single left-finger (GelSight) tactile sensor.
F.5 Representative Rollouts of Visuotactile Policies
Figs. 8 and 9 visualize policy rollouts for two representative tasks from the ManiFeel benchmark: tactile sorting under dim lighting and blind insertion in a vision-occluded environment. These visualizations show the temporal evolution of observations from multiple sensory channels, including front-facing camera view and tactile inputs, across key action stages. Such rollouts help illustrate how policies leverage tactile feedback when visual inputs are ambiguous or occluded. In the tactile sorting task, for example, tactile observations enable the policy to distinguish between objects with similar visual appearances, such as a table tennis ball and a golf ball, by identifying differences in surface texture. This allows the policy to correct initial grasping errors, such as mistakenly picking up the table tennis ball, by regrasping and successfully retrieving the target golf ball to complete the sorting task. This behavior is particularly valuable under dim lighting conditions, where visual cues alone are insufficient for reliable classification. In the blind insertion task, tactile feedback plays a critical role in implicitly estimating the in-hand pose of the peg, which is otherwise unobservable due to the occluded visual field inside the closed container. By adapting the grasp and alignment based on contact signals, the policy is able to achieve successful insertions without relying on external visual confirmation. These examples highlight the importance of tactile sensing for robust policy execution in vision-compromised scenarios. Additional examples and full rollouts can be found in the accompanying supplementary video.
F.6 Real-world Experiment
In the real-world experiment, we evaluate the tactile sorting task, where the robot must classify and sort two visually similar spherical objects, a golf ball and a table tennis ball, into designated bowls. The robot is equipped with a front-facing Intel RealSense camera and a single left-finger tactile sensor (GelSight Mini) for perception. Control is performed using an OSC-yaw controller, which regulates the end-effector’s position and yaw rotation while keeping roll and pitch fixed.
Fig. 10 presents visuotactile policy rollouts under both (a) normal and (b) dim lighting conditions. These visualizations demonstrate how tactile feedback enhances policy robustness, particularly when visual inputs are degraded. For instance, in low-light scenarios, tactile sensing enables the robot to distinguish object surface features and correct initial grasping misalignment. The successful completion of the sorting task under both lighting conditions confirms the value of incorporating tactile perception. Furthermore, the close resemblance between simulated and real tactile feedback, along with consistent policy behavior, highlights the potential of the ManiFeel simulation benchmark and datasets for enabling sim-to-real transfer in visuotactile policy learning. Additional examples and full rollouts can be found in the accompanying supplementary video.

### v2

@@ D-C Runtime Analysis of Tactile Representations
In this section, we evaluate the training time required for different tactile encoders and representations, including ResNet18, UniT, T3, and AnyTouch. Table VIII summarizes the training time per epoch and total training time (over 20 epochs) for visuotactile policies (Vision+TacRGB sensing configuration) in the ball sorting task. All experiments are conducted using an NVIDIA A30 GPU with a batch size of 8. AnyTouch incurs substantially higher computational cost, with each training epoch taking over 67 minutes. In contrast, UniT and T3 are more efficient, with training times of approximately 26 and 18 minutes per epoch, respectively, making them better suited for faster policy iteration and development.
@@
TABLE VIII: Average training time per epoch and total training time (over 20 epochs) for visuotactile policies with different tactile representations in the real-world sorting task under normal lighting conditions. The dataset includes 103 human demonstrations. Time per epoch is reported in minutes and total time in hours.
@@
D-D Experiment on Different Camera Viewpoints
Fig. 13 illustrates the performance of policies in the object-search task when using visual observations from two different camera viewpoints, specifically wrist-mounted and front-facing perspectives. Fig. 13a illustrates tactile inputs and visual images from occluded front views and dim wrist views. The result shows that both vision-only and visuotactile (vision+tacRGB) policies achieve higher success rates when trained with wrist-mounted inputs (see Fig. 13b). Moreover, the visuotactile policy consistently outperforms its vision-only counterpart under both viewpoints, confirming that incorporating tactile sensing enhances performance under visually limited conditions.
@@ D-D Experiment on Different Camera Viewpoints
(b) Success rate
Figure 13: Comparison of front and wrist camera views for vision-only and visuotactile (vision + tacRGB) policies in the object-search task. (a) Sample observations showing tactile inputs and visual images from occluded front views and dim wrist views. (b) Final success rates (%) averaged over the last ten epochs and three seeds.